# Affect, Body, Cognition, Demographics, and Emotion: The ABCDE of Text Features for Computational Affective Science

### v1

@@ Abstract
Work in Computational Affective Science and Computational Social Science explores a wide variety of research questions about people, emotions, behavior, and health. Often they make use of language data that is first labeled with relevant information such as the use of emotion words and age of the speaker. Even though many resources and algorithms exist to enable such labeling, finding and using them is still a substantial impediment, especially to practitioners in fields outside of computer science. Here, we present the ABCDE dataset (“Affect, Body, Cognition, Demographics, and Emotion”), a large-scale collection of over 400 million text utterances from social media, blogs, books, and AI-generated sources, annotated for a number of features relevant to computational affective and social science. ABCDE facilitates inter-disciplinary research in wide range of fields, including affective science, cognitive science, the digital humanities, sociology, political science, and computational linguistics.
@@
wahle@uni-goettingen.de; vkpriya@cs.toronto.edu
@@ 1.  Introduction
In the domain of Computational Affective Science (CAS), research has shown that language use is strongly associated with, and indicative of, mental states and cognitive processes in the human brain Guntuku et al. (2017). When we are sad, we express ourselves using words that are indicative of or associated with low sentiment or valence, or even subtle linguistic signals such as shorter, more terse utterances. An excessive use of qualifiers such as “I think", “probably", and “right?" suggests a lack of confidence or assertiveness in one’s statements. Linguistic properties of utterances are thus immensely useful signals of our mental models and processes, how we as humans perceive ourselves and the world around us.
@@
Our primary motivation in creating this dataset is to unify textual features that encode aspects of behavior and cognition that are of interest to researchers in the affective sciences. While emotion and affect have been prominent topics of study here, recent advances in embodied cognition emphasize how bodily awareness and interactions with the environment shape many of these cognitive functions, and have downstream correlations with mental, emotional, and physical health Wu et al. (2025). We therefore include body part mentions as a core feature of our dataset, allowing researchers to quantify patterns at the intersection of bodily mentions and emotions/affect. With Demographic information, we also consider features that encode social context, in addition to linguistic and lexical features of the text itself. This is a crucial variable for most research in the affective and social sciences, where concepts like emotion and cognition are not constant, stable traits, but demonstrate significant variation with demographic features like age, within and across population sub-groups — in fact, it is this variation that is often the phenomenon of interest.
@@ 2.  Related Work
There exists a small but growing set of accessible text analysis toolkits that have a narrower focus on specific domains of research. The Cornell Conversational Analysis Toolkit Chang et al. (2020) is a collection of datasets and computational scripts that enable research into aspects of social networks, conversational dynamics, and the sociolinguistics of online interactions. GutenTag is an NLP-driven tool for Digital Humanities research in particular, with a web-based interface that automatically extracts several features of interest from literary texts Brooke et al. (2015). In the Affective Sciences domain, packages like VADER Hutto and Gilbert (2014) and Textblob [^2] https://textblob.readthedocs.io/ are widely used for sentiment analysis. The Emotion Dynamics toolkit Vishnubhotla and Mohammad (2022) provides scripts to quantify patterns of change in emotional expression in textual utterances over time.
@@ 3.1.  Data Sources
Books (Google, 1800–2012): 177.1M 5-grams from the English Fiction subset of the Google Books Ngram Corpus (version 20120701) Google Inc. (2012). Each 5-gram is treated as an instance.
@@
AI-Generated Texts (Various, 2022–2025): 68.9M AI-completions from 15 distinct datasets, including conversational texts (WildChat-1M Zhao et al. (2024), LMSYS-Chat-1M Zheng et al. (2024), PIPPA Gosling et al. (2023), HH-RLHF Bai et al. (2022), Prism Kirk et al. (2024), and APT Wahle et al. (2022); persuasive essays (Anthropic-Persuasiveness Durmus et al. (2024)); AI text detection datasets (M4 Wang et al. (2024), MAGE Li et al. (2024), LUAR Soto et al. (2024)); reasoning traces (General Thoughts 430k Reasoning (2024), Reasoning Shield Li et al. (2025), SafeChain Jiang et al. (2025), STAR-1 Wang et al. (2025)); and narratives (TinyStories Eldan and Li (2023)). For conversational datasets, we define the LLM response from a single user–chatbot interaction turn as a text instance (multi-turn conversations are split into multiple instances). Preference datasets generally comprise of a pair of LLM generations (in response to a user prompt) along with an indicator of the preferred generation; we take each generation separately to be an instance of AI-generated text. The AI-generated text detection datasets consist of LLM generations of varying lengths, intended to simulate human text in specific domains like news articles and social media posts; each generation is considered an instance. For reasoning traces, we disregard the final model output and only use the chain-of-thought text for each query as the instance.
@@ 3.2.  Annotated Features
We extract linguistic features through a set of lexicons, word lists, and regular expressions, organized along five dimensions relevant to computational affective science: Affect, Emotion, Body, Cognition, and Demographics. Word lists and lexicons as a measurement method offer several advantages, particularly in inter-disciplinary studies: ease-of-use, generalization power, reliability, adaptability, and interpretability, among others. Variation in word choices across populations and data subsets is also intertwined with multiple cognitive and social processes relating to language use (such as power dynamics, cultural and demographic factors, medium of communication, etc.).
@@
Emotion: We compute average, count, and binary presence flag features for multiple discrete emotions: the eight basic emotions of anger, anticipation, disgust, fear, joy, sadness, surprise, and trust, computed using the NRC Emotion Intensity lexicon Mohammad (2018b); warmth, competence, sociability, and trust using the WCST Lexicon Mohammad (2025); and anxiety and calmness with the NRC WorryWords lexicon Mohammad (2024).
Body: We identify Body Part Mentions (BPMs) by matching text tokens against curated lists of 292 anatomical unigrams, bigrams, and trigrams from Zhuang et al. (2024); Wu et al. (2025), capturing mentions with possessive pronouns (e.g., “my heart”, “her hand”). Features are recorded as binary flags for the presence of a body part mention, and for each possessive pronoun, as a list of BPMs used with that pronoun (i.e, MyBPM is a list containing BPMs used with the pronoun ‘my’).
@@
Cognition: We classify cognitive processes by identifying “thinking words” derived from categories outlined in Bloom’s Taxonomy and Queensland’s Glossary of Cognitive Verbs. [^3] https://adp.uni.edu/documents/bloomverbscognitiveaffectivepsychomotor.pdf, [^4] https://www.qcaa.qld.edu.au/downloads/senior-qce/common/snr_glossary_cognitive_verbs.pdf We categorize 98 unigrams into 11 categories (e.g., analyzing, learning, decision-making). A binary flag feature indicating presence is recorded for each category.
@@
We provide a detailed overview of regexes used in our code repository.
@@
In other words, while lexicon-based methods are not as accurate as neural models at instance-level estimation (i.e, for estimating the sentiment of a particular sentence), they are comparable and sufficient when used as tools of aggregate-level analysis. This means that relative patterns of change, for comparing emotional intensity across different sources, or quantifying the patterns of changes in sentiment over a temporal period (longitudinal analysis), can be estimated with high accuracy using lexicon-based methods.
In Teodorescu and Mohammad (2023), the authors empirically demonstrate that emotion arcs (i.e, temporal fluctuations of valence intensity) obtained with word-level lexicons of valence highly correlate with the ground-truth arcs (with correlation scores >0.9>0.9) for social media texts, provided certain hyper-parameters (like the width of the temporal window) are appropriately set. In the domain of digital humanities, Öhman et al. (2024) measure the agreement of lexicon-generated arcs for classic fiction novels with human annotation, and find high consensus. In their work, the NRC emotion intensity lexicon was customized to the literary domain with a word-similarity measure, which is one of the recommended practices for the use of emotion lexicons Mohammad (2023).
We note that ABCDE can in the future be expanded to annotate emotion and affect features using highly-accurate instance-level estimation models; however, the measurements obtained with lexicon-based features provide the necessary utility for the aggregate-level analysis that is of interest in the social sciences.
@@ 4.  Key Research Questions
Figure 5: Q2. Body: % of possessive BPMs for the top ten most common BPMs, across sources.
Figure 6: Q3. Cognition: % of instances with a positive flag for cognitive terms from different categories.
@@ 5.  Conclusion
In this work, we presented ABCDE , a collection of text datasets and linguistic annotations that capture features of interest at the intersection of language, affective science, cognition, and social science. We make all artifacts of ABCDE available, along with the associated code.
@@ 6.  Bibliographical References
  • Mohammad (2018a) Saif M. Mohammad. 2018a. Obtaining reliable human ratings of valence, arousal, and dominance for 20,000 english words. In Proceedings of The Annual Conference of the Association for Computational Linguistics (ACL), Melbourne, Australia.
@@
  • Mohammad (2023) Saif M. Mohammad. 2023. Best practices in the creation and use of emotion lexicons. In Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics, Dubrovnik, Croatia. Association for Computational Linguistics.
@@
  • Mohammad (2025) Saif M Mohammad. 2025. Words of warmth: Trust and sociability norms for over 26k english words. arXiv preprint arXiv:2506.03993.

### v2

@@ Abstract
Work in Computational Affective Science and Computational Social Science explores a wide variety of research questions about people, emotions, behavior, and health. Often they make use of language data that is first labeled with relevant information such as the use of emotion words and age of the speaker. Even though many resources and algorithms exist to enable such labeling, finding and using them is still a substantial impediment, especially to practitioners in fields outside of computer science. Here, we present the ABCDE dataset (“Affect, Body, Cognition, Demographics, and Emotion”), a large-scale collection of over 400 million released text instances from social media, blogs, books, and AI-generated sources, annotated for a number of features relevant to computational affective and social science. ABCDE facilitates inter-disciplinary research in wide range of fields, including affective science, cognitive science, the digital humanities, sociology, political science, and computational linguistics.
@@
wahle@uni-goettingen.de; vkpriya@cs.toronto.edu, saif.mohammad@nrc-cnrc.gc.ca
@@ 1.  Introduction
In the domain of Computational Affective Science (CAS), research has shown that language use is strongly associated with, and indicative of, mental states and cognitive processes in the human brain Guntuku et al. (2017). When we are sad, we express ourselves using words associated with low valence (negative sentiment), or even subtle linguistic signals such as shorter, more terse utterances. When we lack confidence, we tend use more qualifiers such as “I think", “probably", and “right?". And so on. Linguistic properties of utterances are thus immensely useful signals of our mental models and processes, as well as how we as humans perceive ourselves and the world around us. At the same time, language is only an indirect trace of the underlying states we care about. Many mental states and cognitive processes are latent and must be inferred from behavior and context. Thus while it is difficult to determine the exact mental state of an individual from an utterance (something we do not recommend pursuing automatically anyway), the power of language analysis lies in bringing forth trends and patterns of mental states of whole populations (aggregate-level analysis). Language use remains a natural and scalable behavioral trace that can be used to infer patterns of relative differences and change at an aggregate level. For example, how has the use of emotion words by a population changed with time and what psychological changes that indicates; how a cohort of people uses warmth and competence words when posting about different social groups; what moral frames are invoked when talking about controversial issues (such as immigration), etc.
@@
Our primary motivation in creating this dataset is to unify textual features that encode aspects of behavior and cognition that are of interest to researchers in the affective sciences. While emotion and affect have been prominent topics of study here, recent advances in embodied cognition emphasize how bodily awareness and interactions with the environment shape many of these cognitive functions, and have downstream correlations with mental, emotional, and physical health Wu et al. (2025). We therefore include body part mentions as a core feature of our dataset, allowing researchers to quantify patterns at the intersection of bodily mentions and emotions/affect. With Demographic information, we also consider features that encode social context, in addition to linguistic and lexical features of the text itself. This is a crucial variable for most research in the affective and social sciences, where concepts like emotion and cognition are not constant, stable traits, but demonstrate significant variation with demographic features like age, within and across population sub-groups — in fact, it is this variation that is often the phenomenon of interest Hoemann et al. (2025); Gutchess and Rajaram (2023). Age-related differences have been reported in the arousal structure of emotion concepts Trnka et al. (2022), and age and culture can jointly shift cognitive performance and style Na et al. (2017).
@@ 2.  Related Work
There exists a small but growing set of accessible text analysis toolkits that have a narrower focus on specific domains of research. The Cornell Conversational Analysis Toolkit Chang et al. (2020) is a collection of datasets and computational scripts that enable research into aspects of social networks, conversational dynamics, and the sociolinguistics of online interactions. GutenTag is an NLP-driven tool for Digital Humanities research in particular, with a web-based interface that automatically extracts several features of interest from literary texts Brooke et al. (2015). In the Affective Sciences domain, packages like VADER Hutto and Gilbert (2014) and Textblob [^2] https://textblob.readthedocs.io/ are widely used for sentiment analysis. The Emotion Dynamics toolkit Vishnubhotla and Mohammad (2022) provides scripts to quantify patterns of change in emotional expression in textual utterances over time. [^3] https://github.com/Priya22/EmotionDynamics
@@ 3.1.  Data Sources
Books (Google, 1800–2012): 177.1M 5-gram occurrences (1.74M unique 5-grams) from the English Fiction subset of the Google Books Ngram Corpus (version 20120701) Google Inc. (2012). Each 5-gram is treated as an instance. While these instances are short, they remain useful for large-scale frequency and longitudinal analyses.
@@
AI-Generated Texts (Various, 2022–2025): 68.9M AI-completions from 15 distinct datasets, including conversational texts (WildChat-1M Zhao et al. (2024), LMSYS-Chat-1M Zheng et al. (2024), PIPPA Gosling et al. (2023), HH-RLHF Bai et al. (2022), Prism Kirk et al. (2024), and APT Wahle et al. (2022); persuasive essays (Anthropic-Persuasiveness Durmus et al. (2024)); AI text detection datasets (M4 Wang et al. (2024), MAGE Li et al. (2024), LUAR Soto et al. (2024)); reasoning traces (General Thoughts 430k Reasoning (2024), Reasoning Shield Li et al. (2025), SafeChain Jiang et al. (2025), STAR-1 Wang et al. (2025)); and narratives (TinyStories Eldan and Li (2023)). For conversational datasets, we define the LLM response from a single user–chatbot interaction turn as a text instance (multi-turn conversations are split into multiple instances). Preference datasets generally comprise of a pair of LLM generations (in response to a user prompt) along with an indicator of the preferred generation; we take each generation separately to be an instance of AI-generated text. The AI-generated text detection datasets consist of LLM generations of varying lengths, intended to simulate human text in specific domains like news articles and social media posts; each generation is considered an instance. For reasoning traces, we disregard the final model output and only use the chain-of-thought text for each query as the instance. Instances are multi-feature. One text can activate several feature families, and some short texts may not activate any feature in a given family.
@@ 3.2.  Annotated Features
We extract linguistic features through a set of lexicons, word lists, and regular expressions, organized along five dimensions relevant to computational affective science: Affect, Emotion, Body, Cognition, and Demographics. Word lists and lexicons as a measurement method offer several advantages, particularly in inter-disciplinary studies: ease-of-use, generalization power, reliability, adaptability, and interpretability, among others. Variation in word choices across populations and data subsets is also intertwined with multiple cognitive and social processes relating to language use (such as power dynamics, cultural and demographic factors, medium of communication, etc.). Appendix A.1 summarizes the released resources, and Appendix A.2 reports per-source text statistics.
@@
Emotion: We compute average, count, and binary presence flag features for multiple discrete emotions: the eight basic emotions of anger, anticipation, disgust, fear, joy, sadness, surprise, and trust, computed using the NRC Emotion Intensity lexicon Mohammad (2018b); warmth, competence, sociability, and trust using the WCST Lexicon Mohammad (2025a); and anxiety and calmness with the NRC WorryWords lexicon Mohammad (2024).
Body: We identify Body Part Mentions (BPMs) by matching text tokens against curated lists of 292 anatomical unigrams, bigrams, and trigrams from Zhuang et al. (2024); Wu et al. (2025), capturing mentions with possessive pronouns (e.g., “my heart”, “her hand”). Features are recorded as binary flags for the presence of a body part mention, and for each possessive pronoun, as a list of BPMs used with that pronoun (i.e, MyBPM is a list containing BPMs used with the pronoun ‘my’). Appendix A.3 lists the body part words used.
@@
Cognition: We classify cognitive processes by identifying “thinking words” derived from categories outlined in Bloom’s Taxonomy and Queensland’s Glossary of Cognitive Verbs. [^4] https://adp.uni.edu/documents/bloomverbscognitiveaffectivepsychomotor.pdf, [^5] https://www.qcaa.qld.edu.au/downloads/senior-qce/common/snr_glossary_cognitive_verbs.pdf We categorize 98 unigrams into 11 categories (e.g., analyzing, learning, decision-making). A binary flag feature indicating presence is recorded for each category. Appendix A.4 lists the full grouped inventory.
@@
These regexes are not exhaustive but target high-precision first-person self-disclosure patterns and public controlled vocabularies. Appendix A.5 shows the full list of patterns used.
@@
In other words, while lexicon-based methods are not as accurate as neural models at instance-level estimation (i.e, for estimating the sentiment of a particular sentence), they are comparable and sufficient when used as tools of aggregate-level analysis. This means that relative patterns of change, for comparing emotional intensity across different sources, or quantifying the patterns of changes in sentiment over a temporal period (longitudinal analysis), can be estimated with high accuracy using lexicon-based methods. The source resources also provide validation evidence for aggregate use Mohammad (2018a); Mohammad and Turney (2013); Zhuang et al. (2024).
In Teodorescu and Mohammad (2023), the authors empirically demonstrate that emotion arcs (i.e, temporal fluctuations of valence intensity) with word-level lexicons of valence highly correlate with the ground-truth arcs (with correlation scores >0.9>0.9) for social media texts, provided certain hyper-parameters (like the width of the temporal window) are appropriately set. In the domain of digital humanities, Öhman et al. (2024) measure the agreement of lexicon-generated arcs for classic fiction novels with human annotation, and find high consensus. In their work, the NRC emotion intensity lexicon was customized to the literary domain with a word-similarity measure, which is one of the recommended practices for the use of emotion lexicons Mohammad (2023).
@@ 4.  Key Research Questions
Figure 5: Q2. Body: % of possessive BPMs for the top ten most common BPMs, across sources. Dashed lines represent the average occurence of BPMs per dataset.
Figure 6: Q3. Cognition: % of instances with a positive flag for cognitive terms from different categories. Dashed lines represent the average occurence of cognitive terms per dataset.
@@ 5.  Conclusion
In this work, we presented ABCDE, a collection of text datasets and linguistic annotations that capture features of interest at the intersection of language, affective science, cognition, and social science. We make all artifacts of ABCDE available, along with the associated code. Notably, several recent works already use ABCDE, including sentence-level annotations of social perception dimensions Ayesh et al. (2026), analyses of emotion expression across the lifespan in social media Teodorescu et al. (2026), and analyses for large-scale lexical norms for warmth and trust Mohammad (2025b), and anxiety Mohammad (2026).
@@
Future work will add information for more affective, social perception, and moral dimensions: notably the care–harm, authority–subversion, loyalty–betrayal, fairness–unfairness, and purity–degradation dimensions from the NRC Moral Foundations Lexicon (based on the Moral Foundation Theory of Haidt and Graham (2007)). We also hope that future work led by native speakers of various languages will include datasets in various languages.
@@
  5. 5.
    Our feature set does not model syntax, discourse structure, or compositional effects. Even so, prior lexicon-based studies show that large-scale comparative trends can remain informative at aggregate scale Teodorescu and Mohammad (2023); Öhman et al. (2024).
@@
  4. 4.
    The released instance-level annotations include demographic attributes such as age, gender, religion, occupation, and location. We strictly prohibit any commercial use of our dataset, or its use for profiling, targeting, or individual-level inference.
Acknowledgments
This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Thanks to Terry Ruas and Lars Kaesberg for early feedback on this work.
@@ 6.  Bibliographical References
  • Ayesh et al. (2026) Mutaz Ayesh, Saif M. Mohammad, and Nedjma Ousidhoum. 2026. Annotating dimensions of social perception in text: The first sentence-level dataset of warmth and competence.
@@
  • Covington and McFall (2010) Michael A. Covington and Joe D. McFall. 2010. Cutting the gordian knot: The moving-average type-token ratio (MATTR). Journal of Quantitative Linguistics, 17(2):94–100.
@@
  • Gutchess and Rajaram (2023) Angela Gutchess and Suparna Rajaram. 2023. Consideration of culture in cognition: How we can enrich methodology and theory. Psychonomic Bulletin & Review, 30(3):914–931.
  • Haidt and Graham (2007) Jonathan Haidt and Jesse Graham. 2007. When morality opposes justice: Conservatives have moral intuitions that liberals may not recognize. Social justice research, 20(1):98–116.
  • Hoemann et al. (2025) Katie Hoemann, Yeasle Lee, Èvelyne Dussault, Simon Devylder, Lyle H Ungar, Dirk Geeraerts, and Batja Mesquita. 2025. The construction of emotional meaning in language. Communications Psychology, 3(1):99.
@@
  • Kaesberg et al. (2024) Lars Kaesberg, Terry Ruas, Jan Philip Wahle, and Bela Gipp. 2024. CiteAssist: A system for automated preprint citation and BibTeX generation. In Proceedings of the Fourth Workshop on Scholarly Document Processing (SDP 2024), pages 105–119, Bangkok, Thailand. Association for Computational Linguistics.
@@
  • McCarthy and Jarvis (2010) Philip M. McCarthy and Scott Jarvis. 2010. MTLD, vocd-D, and HD-D: A validation study of sophisticated approaches to lexical diversity assessment. Behavior Research Methods, 42(2):381–392.
  • Mohammad (2018a) Saif M. Mohammad. 2018a. Obtaining reliable human ratings of valence, arousal, and dominance for 20,000 english words. In Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 174–184, Melbourne, Australia. Association for Computational Linguistics.
@@
  • Mohammad (2023) Saif M. Mohammad. 2023. Best practices in the creation and use of emotion lexicons. In Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics, pages 1825–1836, Dubrovnik, Croatia. Association for Computational Linguistics.
@@
  • Mohammad (2025a) Saif M Mohammad. 2025a. Words of warmth: Trust and sociability norms for over 26k english words. arXiv preprint arXiv:2506.03993.
  • Mohammad (2025b) Saif M. Mohammad. 2025b. Words of warmth: Trust and sociability norms for over 26k English words. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 18830–18850, Vienna, Austria. Association for Computational Linguistics.
  • Mohammad (2026) Saif M. Mohammad. 2026. From composure to catastrophe: Norms of calmness–anxiety associations for 54,000 English words and multiword expressions. PsyArXiv preprint.
  • Mohammad and Turney (2013) Saif M. Mohammad and Peter D. Turney. 2013. Crowdsourcing a word-emotion association lexicon. Computational Intelligence, 29(3):436–465.
  • Na et al. (2017) Jinkyung Na, Chih-Mao Huang, and Denise C. Park. 2017. When age and culture interact in an easy and yet cognitively demanding task: Older adults, but not younger adults, showed the expected cultural differences. Frontiers in Psychology, 8:457.
@@
  • Teodorescu et al. (2026) Daniela Teodorescu, Jan Philip Wahle, and Saif M. Mohammad. 2026. Age and affect in language: How emotion expression on social media varies across adulthood. In Proceedings of the 1st Workshop on Computational Affective Science (CAS 2026), Palma de Mallorca, Spain. European Language Resources Association (ELRA).
  • Trnka et al. (2022) Radek Trnka, Josef Mana, and Martin Kuska. 2022. Age-related differences in valence and arousal of emotion concepts. Ageing & Society, 42(9):1991–2007.
@@
  • Wahle et al. (2023) J. Wahle, T. Ruas, S. M. Mohammad, N. Meuschke, and B. Gipp. 2023. Ai usage cards: Responsibly reporting ai-generated content. In 2023 ACM/IEEE Joint Conference on Digital Libraries (JCDL), pages 282–284, Los Alamitos, CA, USA. IEEE Computer Society.